# DynaFlow: Dynamics-embedded Flow Matching for Physically Consistent Motion Generation from State-only Demonstrations

### v2

@@ V-B Real-World Experiments
To validate the motions learned by DynaFlow, we conducted a series of real-world experiments on a Unitree Go1 quadruped robot. For all experiments, we set the integration step size to $\Delta t=1$ to ensure minimal latency for real-time execution on hardware.
As shown in Fig. 4, we first evaluated DynaFlow on trotting and bounding motion learned from the Simulation Rollouts dataset. For deployment, we adopted a receding-horizon control scheme with a 10 Hz replanning rate, executing the first five actions of each new plan. Under this setup, the robot successfully executed locomotion at speeds up to 2.0 m/s for each gait, consistent with the velocity range covered in the training data. Although the training data did not include explicit gait transitions, the robot exhibited smooth and stable switching between gaits at a forward speed of 1.5 m/s. Furthermore, to demonstrate DynaFlow’s ability to produce feasible and coherent long-horizon trajectories, we conducted a challenging open-loop experiment by generating a 450-step (9-second) action trajectory through iterative sampling. Specifically, we repeatedly used the final state of a predicted trajectory segment as the initial state for generating the next segment, sequentially chaining them together. Using this trajectory, the robot successfully traversed a 4 m forward path within a 1 m-wide corridor using both gaits, executing it without any replanning. This highlights DynaFlow’s ability to maintain long-term consistency, enabling entire sequences to run on real hardware without replanning.
We further evaluated DynaFlow under a dynamically infeasible kinematic demonstration using the Retargeted Motion Capture dataset. In this experiment, we used a Vicon motion capture system to obtain the linear velocity expressed in the base frame for observation.
The results highlight DynaFlow’s ability to transform infeasible reference trajectories into physically grounded behaviors while preserving their stylistic characteristics. As shown in Fig. 5, the model first transforms the inconsistent kinematic demonstration into a dynamically feasible trajectory in simulation. The corrected trajectory was executed on the real robot without any fine-tuning or adaptation, resulting in stable galloping behavior.
@@ V-C Robustness to External Disturbances
As illustrated in Fig. 7, DynaFlow consistently achieves a higher or comparable survival rate than the SA-Rollout baseline across most configurations. Furthermore, DynaFlow’s performance remains remarkably consistent across different window sizes and integration timesteps. We attribute this enhanced robustness to the rich learning signal provided by the differentiable simulator. Unlike SA-Rollout, which is trained to imitate action sequences purely, DynaFlow receives gradients through the dynamics model during its training. This process implicitly encourages the model to learn not just the actions to follow a trajectory, but also the corrective actions needed to remain on a stable path, thus improving its resilience to unforeseen perturbations.
@@
VI Conclusion
In this paper, we introduced DynaFlow, a novel generative framework that addresses the challenges of physical inconsistency and action data reliance in robot control. By embedding a differentiable simulator into a flow matching model, DynaFlow guarantees the dynamic feasibility of the generated motions by construction. Its end-to-end differentiable nature enables training on state-only demonstrations, allowing it to infer the actions required to produce a given motion. Our quantitative analysis demonstrated that DynaFlow produces strictly feasible motions even from physically inconsistent data while preserving distributional fidelity. Our real-world deployments further showcased the practical effectiveness of our approach, successfully executing a diverse range of locomotions under both closed-loop and open-loop conditions.

### v3

@@ V-B Real-World Experiments
To validate the motions learned by DynaFlow, we conducted real-world experiments on a Unitree Go1 quadruped robot. For all experiments, we set the integration step size to $\Delta t=1$ to minimize latency for real-time hardware execution.
As shown in Fig. 4, we evaluated DynaFlow on trotting and bounding motions learned from the Simulation Rollouts dataset. For deployment, we adopted a receding-horizon scheme with a 10 Hz replanning rate, executing the first five actions of each plan. Under this setup, the robot successfully executed locomotion at speeds up to 2.0 m/s for each gait, consistent with the velocity range covered in the training data. Although explicit gait transitions were absent from the training data, the robot exhibited smooth and stable switching between gaits at a forward speed of 1.5 m/s. To further assess DynaFlow’s long-horizon feasibility and coherence, we conducted a challenging open-loop experiment by generating a 450-step (9-second) action trajectory through iterative sampling. We repeatedly used the final state of a predicted trajectory segment as the initial state for generating the next segment, sequentially chaining them together. Using this trajectory, the robot successfully traversed a 4 m forward path within a 1 m-wide corridor using both gaits, without any replanning. This highlights DynaFlow’s ability to maintain long-term consistency, enabling entire sequences to run on real hardware in open-loop.
We further evaluated DynaFlow on a dynamically infeasible kinematic demonstration using the Retargeted Motion Capture dataset. In this experiment, a Vicon motion capture system provided base-frame linear velocity for observations.
The results highlight DynaFlow’s ability to transform infeasible reference trajectories into physically grounded behaviors while preserving stylistic characteristics. As shown in Fig. 5, the model transforms the inconsistent kinematic demonstration into a dynamically feasible trajectory in simulation, which is then executed on the real robot without fine-tuning or adaptation, resulting in a stable galloping motion.
@@ V-C Robustness to External Disturbances
As illustrated in Fig. 7, DynaFlow achieves a higher or comparable survival rate than the SA-Rollout baseline across most configurations. Furthermore, DynaFlow’s performance remains remarkably consistent across different window sizes and integration timesteps. We attribute this enhanced robustness to the rich learning signal provided by the differentiable simulator. Unlike SA-Rollout, which directly imitates action sequences, DynaFlow receives gradients through the dynamics model during its training. This process implicitly encourages the model to learn not just the actions to follow a trajectory, but also the corrective actions needed to remain on a stable path, thus improving its resilience to unforeseen perturbations.
@@
VI Discussion
In this paper, we introduced DynaFlow, a novel generative framework that addresses the challenges of physical inconsistency and action data reliance in robot control. By embedding a differentiable simulator into a flow matching model, DynaFlow guarantees dynamically feasible motions by construction. Its differentiability enables end-to-end training on state-only demonstrations, inferring actions to reconstruct a given motion. Quantitative results show that DynaFlow produces strictly feasible motions even from physically inconsistent data while preserving distributional fidelity. Our real-world deployments further showcased its practical effectiveness, successfully executing diverse locomotions under both closed-loop and open-loop conditions.
Future directions include several extensions. First, while our current implementation utilizes a fixed horizon of $H=16$, analyzing gradient stability over extended horizons will be crucial for long-term tasks. Second, expanding to more complex platforms with rich dynamically infeasible datasets, such as humanoids, will offer a more rigorous evaluation of the framework’s robustness. Finally, while this work primarily focuses on locomotion, extending DynaFlow to dynamic loco-manipulation and contact-rich object interaction is a promising direction. Our supplementary video includes a preliminary demonstration of a ball manipulation task.